OpenStudy (v-sync):

Quadrilateral QRST is a square. If the measure of angle RQS is 5x – 15. Find the value of x. give answer and show all work

OpenStudy (v-sync):

@Kevin

OpenStudy (v-sync):

@Kevin

OpenStudy (natalief):

(5x – 15) x 8 = 360

OpenStudy (natalief):

x=12

OpenStudy (natalief):

or you can simply calculate 5x-15 = 45

Hayhayz (hayhayz):

Whats up with the mass tagging

NvidiaIntely (nvidiaintely):

@Awolflover1 @AloneS @Astrophysics @AnonymousHD @awesomedude2078 @aly_cat96 @Bobo-i-bo @beggi9 @Benji @brady05 @countrygirl1431 @callielovesyhuu @civicsiscool44 @chpatterson @caseysorensen @charleskelsey1 @Devonhoward15 @DallasCowboys88 @dolphin1998 @dunk789 @danieldjpon3 @emmigrace222 @Ebayminer126 @Ellie201 @Frozen15 @Foxythepirate0984 @FancySharkk @GracyGirl @GIitter @gnech @ganeshie8 @htxatlarge @hadeel @hhammon4 @Helpmeeasetjx @inkyvoyd @Jumperman @kidrah69 @LunyMoony @Lord2clash @Living4Jesus @MrNood @NatalieF

Hayhayz (hayhayz):

Please stop. Do not tag a bunch of people like that.

NvidiaIntely (nvidiaintely):

Its just that @V-Sync needs help here and no ones coming to help :(

Hayhayz (hayhayz):

@NatalieF did not help?

NvidiaIntely (nvidiaintely):

No, she said she did not understand that .

Hayhayz (hayhayz):

Well only tag a few people such as mods and qualified helpers. Not everyone.

Hayhayz (hayhayz):

Or post the question link in the chats

OpenStudy (natalief):

I have post the answer actually...

OpenStudy (frozen15):

Please don't mass tag , that is against COC , only a few is fine but not over at least 7 people

NvidiaIntely (nvidiaintely):

Okay fine.

Hayhayz (hayhayz):

Thank you

OpenStudy (dallascowboys88):

x=12

OpenStudy (dallascowboys88):

I Agree With @NatalieF

OpenStudy (caseysorensen):

its x=21

NvidiaIntely (nvidiaintely):

BTW what is wrong with mass tagging? Why is it bad to call a bunch of people?

OpenStudy (dallascowboys88):

Mass Tagging Is Considered Spamming,If It's For No Reason Am I Right?

OpenStudy (dallascowboys88):

I'm Pretty Sure It Is.

OpenStudy (frozen15):

thanks @NvidiaIntely we are just saying not to do that because , its against COC for 1 and for 2 , it would be your luck a mod sees this and warns you , they are strict...

OpenStudy (dallascowboys88):

Yes We Are I Agree With @Frozen15

OpenStudy (dallascowboys88):

She Deserves A Medal Don't You Guys Think?

OpenStudy (dallascowboys88):

And Give Me One Too If Youd Like. XP

OpenStudy (v-sync):

My question turned into a chat box about COC :/

OpenStudy (frozen15):

thanks @DallasCowboys88 I don't need a medal cause I'm following the rules or telling someone they should , people tell them all the time and they don't always get rewarded , I don't need a reward for doing something everyone else is already doing

OpenStudy (frozen15):

@V-Sync its not you , we were just saying to not mass tag that's it , nothing else needs to be said

OpenStudy (kevin):

@V-Sync 5x-15 = 90/2 Pls solve the rest by yourself :) I want to to solve it until the end answer but I've always got warning from moderator because of it.

OpenStudy (dallascowboys88):

Yes You Do.For Helping Someone Else/Or A Friend In Need So They Don't Get Suspended/Or In Trouble Bye A Moderator For Spamming And Stuff.

OpenStudy (dallascowboys88):

@Frozen15 I'm Keeping The Medal On Your Post Because,I Think You Deserved It.

OpenStudy (frozen15):

tyvm , I gave you one for helping the asker a bit .. I'm going to leave so I wont spam this post

OpenStudy (kevin):

So, what's your answer @V-Sync ?

OpenStudy (v-sync):

5x - 15 = 90/2 5x - 15 = 45 5x - 15 + 15 = 45 + 15 5x = 60 5x/5 = 60/5 x = 12 @Kevin

OpenStudy (kevin):

it's true

OpenStudy (kevin):

congratulation :D

OpenStudy (v-sync):

Thanks!

OpenStudy (kevin):

yw!